OpenStudy (anonymous):

What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0? slope = -2, y-intercept = 2/3 slope = 2, y-intercept = 2/3 slope = -2, y-intercept = -2/3 slope = 2, y-intercept = -2/3 slope = -2, y-intercept = 3/2

OpenStudy (harsha19111999):

\[3y-6+6x+6-2=0\]

OpenStudy (harsha19111999):

Then you get \[6x+3y-2=0\]

OpenStudy (harsha19111999):

The before equation can be compared with ax+by+c=0

OpenStudy (harsha19111999):

Slope=-a/b =-6/3 = -2

OpenStudy (harsha19111999):

And y intercept = 2/3

OpenStudy (harsha19111999):

You get 1st option

OpenStudy (anonymous):

thank youu

OpenStudy (harsha19111999):

You're welcome